A solution has [H3O+] = 3.0×10−5 M . Use the ion product constant of water
Kw=[H3O+][OH−]
to find the [OH−] of the solution.
Express your answer to two significant figures.

[OH−] = ??? M

A solution has [OH−] = 1.7×10−2 M . Use the ion product constant of water
Kw=[H3O+][OH−]
to find the [H3O+] of the solution.
Express your answer to two significant figures.

[H3O+] = ??? M
1
Expert's answer
2015-11-17T08:53:21-0500
1) The multiplication of concentration of [H3O+] and [OH−] is always equal to Kw. So to find missing concentration we just have to divide Kw to the known concentration. In this case we can estimate the concentration of [OH−] by these formulas:
Kw = [H3O+]*[OH−]
[OH−] = Kw/[H3O+]
Kw = 1×10−14
[OH−] = 1×10−14/3.0×10−5 = 3.3×10−10 M
2) The multiplication of concentration of [H3O+] and [OH−] is always equal to Kw. So to find missing concentration we just have to divide Kw to the known concentration. In this case we can estimate the concentration of [H3O+] by these formulas:
Kw = [H3O+]*[OH−]
[H3O+] = Kw/[OH−]
Kw = 1×10−14
[H3O+] = 1×10−14/1.7×10−2 = 5.9×10−13 M
Answer:
1) [OH−] = 3.3×10−10 M
2) [H3O+] = 5.9×10−13 M